OpenStudy (anonymous):

hello i dont understand what they are asking in this question, "Find all of the solutions that satisfy sin 2x + 3cos, x = 0"? . its a calculus question btw

OpenStudy (anonymous):

it's asking you to solve for x in the equation sin(2x) + 3cos(x) = 0

OpenStudy (anonymous):

could u plz give me few first steps??

OpenStudy (anonymous):

are you know the rule of sin2x

OpenStudy (anonymous):

2sinxcosx??

OpenStudy (anonymous):

yeah apply it then what you get?

OpenStudy (anonymous):

cosx(2sinx+3)??

OpenStudy (anonymous):

right :) then you have to divide all the two sides of the equation by cosx the result is...?

OpenStudy (anonymous):

sinx=-3/2?

OpenStudy (anonymous):

good :) then x=?

OpenStudy (anonymous):

yea i dont understand the next step could u tell me ?

OpenStudy (anonymous):

if you have the equation sintheta=A theta=sin^-1 A

OpenStudy (anonymous):

\[x=\sin ^{-1}\frac{ -3 }{ 2 }\]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

oh ok so do i put it in my calculator??

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

wait i get a math error?? beacuse of that negative sign

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what should i do?

OpenStudy (anonymous):

240 in dgree

myininaya (myininaya):

\[2\sin(x)\cos(x)+3\cos(x)=\cos(x)(2\sin(x)+3)=0\] find when both factor are equal to zero 2sin(x)+3 is never zero because sin(x) is never -3/2 because the range of sin(x) is from -1 to 1 so you only need to solve cos(x)=0

OpenStudy (anonymous):

wait so after that i just look atthe unit circle?

OpenStudy (anonymous):

so the answer is 1 ?

myininaya (myininaya):

nope x isn'1 cos(1) is not 0

myininaya (myininaya):

find for what angles cos is 0

myininaya (myininaya):

when are the x-coordinates on your unit circle sheet 0? for what angles?

OpenStudy (anonymous):

90?

myininaya (myininaya):

that is one yep are there any others?

OpenStudy (anonymous):

270?

myininaya (myininaya):

so it would be \[90^o+180^o \cdot n \text{ where n is an integer }\]

OpenStudy (anonymous):

wait what so the soltion is not 90 and 270??

myininaya (myininaya):

Those are some of the solutions You were asked to find all so I helped you out a bit in doing so. The solution I gave contains both of those solutions and many more.